# www.entalus.com

### v1

@@ 2 Preliminaries
$d(x,y)$ $\leq\mathrm{max}({d(x,y)},{d(y,z)})\mbox{\@.}$ (3)
Example 2 (Discrete Metric)
The discrete metric $d(x,y)$, which is $1$ for $x\neq y$ and 00 otherwise, is ultrametric.
As a consequence of (3) the isosceles triangle principle
$d(x,z)$ $=\mathrm{max}({d(x,y)},{d(y,z)})$ (4)
holds whenever $d(x,y)\neq d(y,z)$. Further immediate consequences of the strong triangle inequality are: (1) every point inside a ball is its center, that is, if $d(x,y)<r$ then $\mathrm{B}({x},{r})=\mathrm{B}({y},{r})$, (2) all balls of strictly positive radius $r$ are clopen, that is both open and closed, (3) if two balls are not disjoint then one is included in the other, (4) the distance $d(B_{1},B_{2})\,\coloneqq\,\inf_{x\in B_{1},y\in B_{2}}d(x,y)$ of two disjoint nonempty balls $B_{1}$,$B_{2}$ is obtained as the distance of two arbitrarily chosen elements $x\in B_{1}$, $y\in B_{2}$. (4) ultrametric spaces are totally disconnected, that is, every superset of a singleton set is disconnected, and (5) a sequence $(x_{k})_{k\in\mathbb{N}}$ in an ultrametric space is Cauchy if and only if $\lim_{k\to\infty}d(x_{k+1},x_{k})=0$. We will also make use of a generalized strong triangle inequality (see Appendix 0.B) for ultrametric distances.
$d(a,C)$ $\leq\max{(d(a,b),d(b,C))}\mbox{\@,}$ (5)
for all $a,b\in M$ and $\emptyset\neq C\subseteq M$. For a proof of the following central fact see also Lemma 16 in the Appendix.
Proposition 1
An ultrametric space $(M,d)$ is spherically complete if $\,\cap\,_{{B\in\cal C}}B\neq\emptyset$ for every chain ${\cal C}$ of balls $B_{0}\supseteq B_{1}\supseteq B_{2}\supseteq\ldots$.
Example 3
Clearly, spherical completeness implies Cauchy completeness.
Proposition 2 (([17], p. 59)
If $(M,d)$ is ultrametric then $({\mathrm{Comp}({M})},\mathcal{H}_{d})$ is Cauchy complete.
For a spherically complete ultrametric space $M$, every strictly contracting $T:M\to M$ has a unique fixpoint in $M$ ([43], Theorem 1). Moreover, if $T:M\to M$ is nonexpansive, that is $d(T(x),T(y))\leq d(x,y)$ for all $x,y\in M$, then either $T$ has at least one fixpoint or there exists a ball $B$ of radius $r>0$ such that $T:B\to B$ and for which $d(u,T(u))=r$ for each $u\in B$  ([43], Theorem 2). Such a ball $B$ is sometimes also said to be minimal $T$-invariant.
Example 4
For the spherically complete ultrametric space $(M,d)$ from Example 3 define $T:M\to M$ such that $\alpha\mapsto\gamma$, $\beta\mapsto\delta$, $\gamma\mapsto\alpha$, and $\delta\mapsto\beta$. Then $T$ has no fixpoint and for each $x\in X$ there exists $r>0$ such that $d(u,T(u))=r$ for all $u\in\mathrm{B}[{x},{r}]$.
3 Streams
An $A$-valued stream in ${A}^{\omega}\,\coloneqq\,\mathbb{N}\to A$, for a given nonempty set $A$ of values, is an infinite sequence $(a_{k})_{k\in\mathbb{N}}$ with $a_{k}\in A$. Depending on the application context, streams are also referred to as discrete streams or signals, $\omega$-streams, $\omega$-sequences, or $\omega$-words. The generating function [12] of a stream $(a_{k})_{k\in\mathbb{N}}$ is a formal power series
$\sum_{k\in\mathbb{N}}a_{k}X^{k}$
$(\sum_{k\in\mathbb{N}}a_{k}X^{k})+(\sum_{k\in\mathbb{N}}b_{k}X^{k})$ $\,\coloneqq\,\sum_{k\in\mathbb{N}}(a_{k}+b_{k})X^{k}$ (6)
$(\sum_{k\in\mathbb{N}}a_{k}X^{k})\cdot(\sum_{k\in\mathbb{N}}b_{k}X^{k})$ $\,\coloneqq\,\sum_{k\in\mathbb{N}}(\sum_{i=0}^{k}a_{i}b_{k-i})X^{k}$ (7)
If $A$ is a field, then ${A}^{\omega}\simeq{A}\llbracket{X}\rrbracket$ is a principal ideal domain with the ideal $(X)$, which consists of all streams $f\in X\cdot{A}^{\omega}$, the only non-zero maximal ideal. Moreover, for $A$ a field or a division ring, $({A}^{\omega},+,(a\cdot)_{a\in K})$ is a linear space, whereby the dot product $(a\cdot f)$, for $f\in{A}^{\omega}$, is defined by $\bar{a}\cdot f$. Unless stated otherwise, we assume the values $A$ in ${A}^{\omega}$ to be a field.
Example 5 (Stream Calculus)
A stream is said to be rational [55] if it can be expressed as a quotient $\nicefrac{{p}}{{q}}$ of polynomial streams $p,q\in{A}[{X}]$ such that $[X^{0}]g\neq 0$.
/1(1 - X) $=(1,1,1,1,\ldots)$
/1(1 - X)2 $=(1,2,3,4,\ldots)$
/1(1 - aX) $=(1,a,a^{2},a^{3}\ldots)$ for $\text{for }a\in A$
/X(1 - X - X2) $=(0,1,1,2,3,5,8,\ldots)$ (Fibonacci)
/2(1 + 1 - 4X) $=(1,1,2,5,14,\ldots)$ (Catalan)
The square root of a stream is defined, for instance, in [52]. Ultimately periodic streams such as $\omega$-words are a special case of rational streams [55]. On the other hand, by Theorem 5.4 in [55], there is no such quotient of polynomials for characterizing the stream $(1,1,0,1,0,0,1,0,0,0,1,0,0,0,0,\ldots)$.
Remark 1 (Rational Streams vs. Rational Functions)
The set of rational streams as defined in Example 5 is a subset of the real-valued formal power series. The ring of rational streams should not be confused, however, with the field ${\mathbb{R}}({X})$ of so-called rational functions with real-valued coefficients, which is the quotient field of the polynomial ring ${\mathbb{R}}[{X}]$. For example, $\nicefrac{{1}}{{X}}\in{\mathbb{R}}({X})$ but $\nicefrac{{1}}{{X}}$ is, by definition, not a rational stream.
The index $v({f})$ of a stream $f\in{A}^{\omega}$ is the minimal $k\in\mathbb{N}$ such that $[X^{k}]f\neq 0$, if any exists; otherwise $v({f})\,\coloneqq\,\infty$.
Proposition 3 (Valuation)
For all $f,g\in{A}^{\omega}$:
  1. 1.
    $v({f})=\infty$ iff $f=0$
  2. 2.
    $v({f\cdot g})=v({f})+v({g})$
  3. 3.
    $v({f+g})\geq\mathrm{min}({v({f})},{v({g})})$
Immediate consequences include: (1) $v({1})=0$, (2) $v({f^{-1}})=-v({f})$ for all streams $f$ for which $f^{-1}$ exists, and, (3) $v({f+g})=\mathrm{min}({v({f})},{v({g})})$ whenever $v({f})\neq v({g})$. By construction,
$|{f}|\,\coloneqq\,2^{-v({f})}\mbox{\@,}$ (8)
[…]
Proposition 4 (Non-Archimedean Absolute Value)
For all $f,g\in{A}^{\omega}$:
  1. 1.
    $0\leq|{f}|\leq 1$
  2. 2.
    $|{f}|=0$ if and only if $f=0$
  3. 3.
    $|{f\cdot g}|=|{f}|\cdot|{g}|$
  4. 4.
    $|{f+g}|\leq\mathrm{max}({|{f}|},{|{g}|})$
Some immediate consequences of Proposition 4 include: (1) $|{1}|=1$, (2) $|{r}|=1$ for all roots $r$ of unity , (3) $|{f^{-1}}|=|{f}|^{-1}$ for streams $f$ with an inverse, (4) $||f|-|g||\leq|f-g|$, and (5) $|f+g|=\mathrm{max}({|f|},{|g|})$ whenever $|f|\neq|g|$. Notice also that stream valuation trivially has the non-Archimedean property $|{1+\cdots+1}|\leq|{1}|=1$.
@@ 3 Streams
the set ${A}^{\omega}$ is a metric with a discrete set of values $\Delta_{d}\,\coloneqq\,\{{2^{-n}}\,\mid\,{n\in\mathbb{N}\cup\{\infty\}}\}$. In fact, $d$ is an ultrametric, since the strong triangle inequality
$d(f,h)$ $\leq max(d(f,g),d(g,h))\mbox{\@.}$ (10)
holds for all streams $f$, $g$, $h$ (see Appendix 0.A). As a consequence, the isosceles triangle principle holds for $d$.
Remark 2
${\mathbb{B}}^{\omega}$ with the ultrametric $d$ is known as the Cantor space, and ${\mathbb{N}}^{\omega}$ is the Baire space.
Proposition 5
@@
Both addition and multiplication of streams are continuous with respect to the topology induced by $d$.
@@
Since this metric-induced topology is identical to the product topology $A^{\mathbb{N}}$, where each copy of $A$ is the discrete topology, Tychonoff’s theorem applies.
Proposition 6
${A}^{\omega}$ is compact if and only if $A$ is finite.
A sequence $(f_{k})_{k\in\mathbb{N}}$ is convergent in ${A}\llbracket{X}\rrbracket$ provided that for every $n\in\mathbb{N}$ there exists $K_{n}\in\mathbb{N}$ and $A_{n}\in A$ such that if $k\geq K_{n}$ then $[X^{n}]f_{k}=A_{n}$. This condition says that if we focus attention on only the $n$–th power of $X$, and consider the sequence $[X^{n}]f_{k}$ of coefficients of $X^{n}$ in $f_{k}$ as $k\to\infty$, then this sequence (of elements of $A$) is eventually constant, with the ultimate value $A_{n}$. In this case,
$f$ $\,\coloneqq\,\sum_{n=0}^{\infty}A_{n}X^{n}$ (11)
is a well-defined formal power series, called the limit of the convergent sequence $(f_{k})_{k\in\mathbb{N}}$. We use the notation $\lim_{k\to\infty}f_{k}=f$ to denote this relationship.
Example 6
  1. 1.
    $\lim_{k\to\infty}X^{k}=0$,
  2. 2.
    $\lim_{k\to\infty}\sum_{i=0}^{k}X^{i}=\nicefrac{{1}}{{(1-X)}}$.
Lemma 1
The ultrametric space $({A}^{\omega},d)$ of streams is spherically complete.
Proof
Let $B_{0}\supseteq B_{1}\supseteq\ldots$ be any non-increasing sequence of non-empty balls in ${A}^{\omega}$. Then, the sequence $(r_{k})_{k\in\mathbb{N}}$ of radii $r_{k}$ of the balls $B_{k}$ is a non-increasing sequence of numbers in $\mathbb{R}$, which, by the discreteness of $d$, either becomes constant or converges to zero. If it becomes constant then $\,\cap\,_{k\in\mathbb{N}}B_{k}$ even contains a ball, and if it converges to zero then this intersection of balls also is nonempty, since, for the ultrametricity of $d$, every point inside a ball is its center.
Lemma 1 implies Cauchy completeness of the metric space of streams (see also Appendix 0.A, Lemma 15 for a direct proof).
Corollary 1
$({A}^{\omega},d)$ is Cauchy complete.
@@
Indeed, ${A}^{\omega}\simeq{A}\llbracket{X}\rrbracket$ is the Cauchy completion of the polynomials ${A}[{X}]$ for the prefix metric $d$. Since any metric space is compact if and only if it is Cauchy complete and totally bounded, Corollary 1 and Proposition 6 together imply that
Corollary 2
$({A}^{\omega},d)$ is totally bounded if and only if $A$ is finite.
Remark 3
The metric subspace ${\mathbb{R}}({X})$ of ${\mathbb{R}}^{\omega}$ is not Cauchy complete. Indeed the sequence $(\sum_{k=0}^{n}\frac{1}{k!}X^{k})_{n\in\mathbb{N}}$ of polynomials is Cauchy but it does not converge in ${\mathbb{R}}({X})$.
Remark 4
The set ${A}(\!({X})\!)$ of formal Laurent series of the form $\sum_{k\geq k_{0}}a_{k}X^{k}$, for $k\in\mathbb{Z}$, is the quotient field of the formal power series ${A}\llbracket{X}\rrbracket$. Now, ${A}(\!({X})\!)$ is spherically complete and therefore also Cauchy complete [16]. Indeed, ${A}(\!({X})\!)$ is the Cauchy completion of ${A}({X})$ for the prefix metric $d$.
Stream Products.
@@
Let ${\cal I}$ be a nonempty index set, $A_{\iota}$ a set of values for each index $\iota\in{\cal I}$, and $(\Pi{\iota\in{\cal I}})\,{{A_{\iota}}^{\omega}}$ the set of ${\cal I}$-indexed product of formal power series. The valuation
$|\bar{f}|_{{\cal I}}$ $\,\coloneqq\,\sup_{\iota\in{\cal I}}|f_{\iota}|$ (12)
@@
for products $\bar{f}$ of the form $(f_{\iota})_{\iota\in{\cal I}}$ is just the supremum of the valuation of its components.
Remark 5
@@
In particular, the valuation of finite dimensional products of the form $(f_{1},\ldots,f_{n})$, for $n\in\mathbb{N}$, is $2^{-k}$, where $k\in\mathbb{N}\,\cup\,\{\infty\}$ is the maximal position such that all prefixes ${f_{i}}\!\!\restriction_{k}$, for $i=1,\ldots,n$, only contain zeros.
Remark 6
Likewise, when interpreting $\bar{f}\in(\Pi{\iota\in{\cal I}})\,{{A_{\iota}}^{\omega}}$ as a (dependent) function with domain ${\cal I}$ and codomains ${A_{\iota}}^{\omega}$ for $\iota\in{\cal I}$, then the valuation $|\bar{f}|_{\cal I}$ is obtained as $\sup_{g\in\bar{f}({\cal I})}|g|$.
@@
A metric on the ${\cal I}$-indexed product space $(\Pi{\iota\in{\cal I}})\,{{A_{\iota}}^{\omega}}$ is induced by the valuation (12).
$d_{\cal I}(\bar{f},\bar{g})\,\coloneqq\,|\bar{f}-\bar{g}|_{\cal I}$ (13)
Lemma 2
@@
For nonempty ${\cal I}$ and $A_{\iota}$ for each $\iota\in{\cal I}$, the space $((\Pi{\iota\in{\cal I}})\,{{A_{\iota}}^{\omega},d_{{\cal I}}})$ is ultrametric and spherically complete.
@@
Since $d_{{\cal I}}$ specializes for a singleton index set to the distance $d$ on streams, we usually drop the subindex ${\cal I}$.
4 Transformers
@@
Stream transformers are the basic building blocks for modeling (discrete) dynamical systems, say $T$, with a vector ${x}$ of $n$ input streams and a vector ${y}$ of $m$ output streams (see Figure 1). We write ${{A}^{\omega,{n}}}$ for the set $({A}^{\omega})^{n}$ of $n$-dimensional vectors of $A$-valued streams.
{tikzpicture}
Figure 1: Stream Transformer.
Definition 1 (Stream Transformers)
@@
A stream transformer is a vector- and multi-valued map $T:{{A}^{\omega,{n}}}\to\mathcal{P}({{{B}^{\omega,{m}}}})$, for $n,m\geq 1$. If $A$ ($B$) is a singleton set, then $T$ is a source (sink). If $|T(f)|=1$ (componentwise) for all $f\in({{A}^{\omega,{n}}})$ then $T$ is a deterministic stream transformer; otherwise the stream transformer is said to be non-deterministic. For a deterministic stream transformer we also write $T:{{{A}^{\omega,{n}}}}\to{{{B}^{\omega,{m}}}}$. A stream transformer is said to be identically empty if $T({{{A}^{\omega,{n}}}})=\emptyset$.
@@
Hereby, the direct image of a stream transformer $T$ with respect to a set $F$ of streams is denoted by $T(F)$.
Remark 7
@@
The restriction to vector-valued stream transformer is mainly motivated by notational convenience, as most of the developments here generalize to transformers with heterogenous coefficient sets of the form
@@
and also to infinite products.
Remark 8
Vector-valued stream transformers ${{{A}^{\omega,{n}}}}\to\mathcal{P}({{{B}^{\omega,{m}}}})$ may be identified with stream transformers in ${(A^{n})}^{\omega}\to\mathcal{P}({(B^{m})}^{\omega})$, because of the one-to-one relationship of ${(A^{n})}^{\omega}$ with ${{A}^{\omega,{n}}}$, and ${(B^{m})}^{\omega}$ with ${{B}^{\omega,{m}}}$. If $A$ ($B$) is a field then $A^{n}$ ($B^{m}$) can be made into a field in the usual way.
Definition 2
Rational stream transformers $T$ map a vector $z$ of $n$ rational input streams to a vector of $m$ rational output streams by means of the linear mapping
$z$ $\mapsto R\cdot z\mbox{\@,}$ (14)
where $R$ is an $(n\times m)$-dimensional matrix $R$ of rational streams.
Example 7
Stream circuits [55] are clocked, hardware-like finite structures, which are obtained by (finite) compositions of the rational stream transformers (for $r\in\mathbb{R}$ and $\tau$ the vector transpose)
@@
for multiplying a stream by a constant, adding two streams, copying, and delaying a stream. Such a stream circuit is visualized graphically in Figure 2.
{tikzpicture}
Figure 2: Finite Stream Circuit.
Remark 9
@@
for the output stream $y$ to obtain $y=(\nicefrac{{1}}{{1-X}})\cdot z$. Thus, the rational stream transformer ${\mathbb{R}}^{\omega}\to{\mathbb{R}}^{\omega}$ that is implemented by the feedback circuit in Figure 2 is given by $z\mapsto(\nicefrac{{1}}{{1-X}})\cdot z$. Since $\nicefrac{{1}}{{1-X}}=(1,1,1,\ldots)$, the output stream $y$ is, by stream multiplication (7), of the form $(\sum_{k=0}^{n}[X^{k}]z)_{n\in\mathbb{N}}$.
Every rational stream transformers can be realized by a stream circuit [55]. In addition, rational stream transformers and weighted automata are equally expressive [55]. Not every feedback loop however gives rise to a stream transformer.
Remark 10
Consider for instance the circuit obtained by replacing the one unit delay $D_{1}$ in the feedback loop in Figure 2 by a $1$-multiplier $M_{1}$. From the definitions of the basic stream transformers we read off the corresponding system of stream equations $h_{1}=1\cdot h_{2}$, $h_{3}=z+h_{1}$, $h_{2}=h_{3}$, and $y=h_{3}$. This implies $h_{3}=z\cdot h_{3}$ and therefore $z=0$. But $z$ is supposed to be an arbitrary input stream. Circuits in which every feedback loop passes through at least one register, however, can be shown to represent a rational stream transformer [55].
Example 8 (Functional Composition)
A formal power series may also be viewed as a stream transformer $f(X)$, where $X$ is now instantiated with an input stream, say, $g(X)$ by the ”composition” $f(g(X))$. Formally, the functional composition $g\,\circ\,f$ of two power series $f\,\coloneqq\,\sum_{k\geq 0}a_{k}X^{k}$ and $g\,\coloneqq\,\sum_{k\geq 1}b_{k}X^{k}$, that is, $b_{0}=0$, is a power series with coefficients
$c_{k}$ $\,\coloneqq\,\sum_{n\geq 0}\sum_{j_{1}+\ldots+j_{n}=k}b_{n}a_{j_{1}}\ldots a_{j_{n}}\mbox{\@.}$ (15)
In case $A$ is a field of characteristic 00, Faà di Bruno’s formula provides a more explicit description of these coefficients.
Example 9 (Compositional Inverse)
Inverses for the functional composition (15) of two power series exist under certain conditions, and may be used for reverse computation. Let $f\in{A}^{\omega}$ with $[X^{0}]f=0$ and $[X^{1}]f$ is an invertible element of $A$. Then there exists the composition inverse $g\in{A}^{\omega}$ with $f\,\circ\,g=1$. In the case when the coefficient ring is a field of characteristic 00, the $k$-th coefficient of $g^{n}$, for $n\in\mathbb{N}^{+}$, is obtained from Lagrange inversion $k[X^{k}]g^{n}=n[X^{-n}]f^{-k}$.
Composing Stream Transformers.
@@
Definition 3
@@
$S;T$ is sequential composition, $S\otimes T$ is parallel composition, $S\mathrm{\sqcup}T$ realizes angelical and $S\mathrm{\sqcap}T$ demonic non-determinism, which is possibly unbounded.
Notice that these definitions work for arbitrary vector- and multi-valued maps.
Example 10
Parallel composition $S\otimes T$ is visualized in Figure 4, where $C:{A}^{\omega}\to{{A}^{\omega,{2}}}$ provides two copies of the input stream, and $M:{{A}^{\omega,{2}}}\to{(A^{2})}^{\omega}$ merges its two input stream $(a_{k})_{k\in\mathbb{N}}$ and $(b_{k})_{k\in\mathbb{N}}$ into the single output stream $((a_{k},b_{k}))_{k\in\mathbb{N}}$.
Example 11
Let $b\in{\mathbb{B}}^{\omega}$, and $f,g\in{A}^{\omega}$; then the conditional stream $(\mathrm{{b}\,?\,{f}\,:\,{g}})$ is the unique stream satisfying the identities
$\mathrm{hd}({\mathrm{{b}\,?\,{f}\,:\,{g}}})$ =$=\begin{cases}\mathrm{hd}({f})&\text{if }\mathrm{hd}({b})=1\\ \mathrm{hd}({g})&\text{if }\mathrm{hd}({b})=0\\ \end{cases}$
$\mathrm{tl}({\mathrm{{b}\,?\,{f}\,:\,{g}}})$ $=(\mathrm{{\mathrm{tl}({b})}\,?\,{\mathrm{tl}({f})}\,:\,{\mathrm{tl}({g})}}\mbox{\@.}$
As a coinductive definition [54], the conditional in Definition 11 is well-defined.
Remark 11
Coinductive stream specifications are also said to be behavioral differential equations by identifying the tail operation $\mathrm{tl}({.})$ with differentiation [53]. These behavioural differential equations have an operational reading, from which algorithms for producing the elements of this infinite stream one-by-one can be easily and progressively derived.
There are unique operators on streams satisfying co-inductive identities for addition, multiplication, and inverse ([53], Theorem 3.1).
Example 12
$\mathrm{tl}({f+g})$ $=$ $\mathrm{tl}({f})+\mathrm{tl}({g})$ $\mathrm{tl}({f\cdot g})$ $=$ $\mathrm{tl}({f})\cdot g+\mathrm{hd}({f})\cdot\mathrm{tl}({g})$ $\mathrm{tl}({f^{-1}})$ $=$ $-(\mathrm{hd}({f^{-1}})\cdot\mathrm{tl}({f}))\cdot f^{-1}$ $\mathrm{hd}({f+g})$ $=$ $\mathrm{hd}({f})+\mathrm{hd}({g})$ $\mathrm{hd}({f\cdot g})$ $=$ $\mathrm{hd}({f})\cdot\mathrm{hd}({g})$ $\mathrm{hd}({f^{-1}})$ $=$ $\mathrm{hd}({f})^{-1}$
Example 13 (Map Transformer)
For $\phi:A\to B$, the stream transformer $\mathrm{M}_{\phi}:{A}^{\omega}\to{B}^{\omega}$ is co-inductively defined as the unique solution of the identities ($f\in{A}^{\omega}$)
$\mathrm{hd}({\mathrm{M}_{\phi}(f)})$ $=\phi(\mathrm{hd}({f}))$
$\mathrm{tl}({\mathrm{M}_{\phi}(f)})$ $=\mathrm{M}_{\phi}(\mathrm{tl}({f}))\mbox{\@.}$
Example 14 (Boolean Transformers)
[…]
${b}\,|\,{c}$ $\,\coloneqq\,\mathrm{{b}\,?\,{1}\,:\,{c}}\mbox{\@.}$
4.1 Refinement
@@
$S\preceq T$ $\,\coloneqq\,(\forall{f\in{A}^{\omega}})\,{S(f)\subseteq T(f)}\mbox{\@,}$ (16)

### v2

@@ 3. Streams
the set ${A}^{\omega}$ is a metric with a discrete set of values $\Delta_{d}\,\coloneqq\,\{{2^{-n}}\,\mid\,{n\in\mathbb{N}\cup\{\infty\}}\}$. In fact, $d$ is an ultrametric, since, by construction, the strong triangle inequality
(11) $d(f,h)$ $\leq max(d(f,g),d(g,h))\mbox{\@.}$
holds for all streams $f$, $g$, $h$.
Proposition 3.1.
@@
Both addition (7) and multiplication (8) of streams are continuous with respect to the topology induced by the prefix metric $d$.
@@
The topology induced by the prefix metric $d$ is identical to the product topology $A^{\mathbb{N}}$, where each copy of $A$ is the discrete topology. Therefore Tychonoff’s theorem applies, and ${A}^{\omega}$ is compact if and only if $A$ is finite.
The following fact can easily be checked from the definition of spherically completeness, which requires that the intersection of non-increasing sequences of non-empty balls in ${A}^{\omega}$ is nonempty.
Lemma 3.2.
$({A}^{\omega},d)$ of streams is spherically complete.
@@
As a consequence, $({A}^{\omega},d)$ is also Cauchy complete, and, indeed, ${A}^{\omega}\simeq{A}\llbracket{X}\rrbracket$ is the Cauchy completion of the polynomials ${A}[{X}]$ for the prefix metric $d$.
@@
Let ${\mathcal{I}}$ be a nonempty index set, $A_{\iota}$ a set of values for each index $\iota\in{\mathcal{I}}$, and $(\Pi{\iota\in{\mathcal{I}}})\,{{A_{\iota}}^{\omega}}$ the set of ${\mathcal{I}}$-indexed product of formal power series. The valuation
(12) $|\bar{f}|_{{\mathcal{I}}}$ $\,\coloneqq\,\sup_{\iota\in{\mathcal{I}}}|f_{\iota}|$
@@
for products $\bar{f}$ of the form $(f_{\iota})_{\iota\in{\mathcal{I}}}$ is just the supremum of the valuation of its components.
@@
In particular, the valuation of finite dimensional products of the form $(f_{1},\ldots,f_{n})$, for $n\in\mathbb{N}$, is $2^{-k}$, where $k\in\mathbb{N}\,\cup\,\{\infty\}$ is the maximal position such that all prefixes ${f_{i}}\!\!\restriction_{k}$, for $i=1,\ldots,n$, only contain zeros. Likewise, when interpreting $\bar{f}\in(\Pi{\iota\in{\mathcal{I}}})\,{{A_{\iota}}^{\omega}}$ as a (dependent) function with domain ${\mathcal{I}}$ and codomains ${A_{\iota}}^{\omega}$ for $\iota\in{\mathcal{I}}$, then the valuation $|\bar{f}|_{\mathcal{I}}$ is obtained as $\sup_{g\in\bar{f}({\mathcal{I}})}|g|$.
@@
A metric on the ${\mathcal{I}}$-indexed product space $(\Pi{\iota\in{\mathcal{I}}})\,{{A_{\iota}}^{\omega}}$ is induced by the valuation (12).
(13) $d_{\mathcal{I}}(\bar{f},\bar{g})\,\coloneqq\,|\bar{f}-\bar{g}|_{\mathcal{I}}$
Lemma 3.3.
@@
For nonempty ${\mathcal{I}}$ and $A_{\iota}$ for each $\iota\in{\mathcal{I}}$, the space $((\Pi{\iota\in{\mathcal{I}}})\,{{A_{\iota}}^{\omega},d_{{\mathcal{I}}}})$ is ultrametric and spherically complete.
@@
Since $d_{{\mathcal{I}}}$ specializes for a singleton index set to the distance $d$ on streams, we usually drop the subindex ${\mathcal{I}}$.
4. Transformers
@@
Stream transformers are the basic building blocks for modeling (discrete) dynamical systems, say $T$, with a vector ${x}$ of $n$ input streams and a vector ${y}$ of $m$ output streams. We write ${{A}^{\omega,{n}}}$ for the set $({A}^{\omega})^{n}$ of $n$-dimensional vectors of $A$-valued streams.
Definition 4.1 (Stream Transformers).
@@
A stream transformer is a vector- and multivalued map $T:{{A}^{\omega,{n}}}\to\mathcal{P}({{{B}^{\omega,{m}}}})$, where $n,m\geq 1$. If $|T(f)|=1$ (componentwise) for all $f\in({{A}^{\omega,{n}}})$ then $T$ is a deterministic stream transformer; otherwise the stream transformer is said to be nondeterministic. For a deterministic stream transformer we also write $T:{{{A}^{\omega,{n}}}}\to{{{B}^{\omega,{m}}}}$.
@@
The direct image of a stream transformer $T$ with respect to a set $F$ of streams is denoted by $T(F)$.
@@
The restriction to vector-valued stream transformers is mainly motivated by notational convenience, as most of the developments in this paper generalize to transformers with heterogenous coefficient sets of the form
@@
and also to infinite products. Vector-valued stream transformers ${{{A}^{\omega,{n}}}}\to\mathcal{P}({{{B}^{\omega,{m}}}})$ may be identified with stream transformers in ${(A^{n})}^{\omega}\to\mathcal{P}({(B^{m})}^{\omega})$, because of the one-to-one relationship of ${(A^{n})}^{\omega}$ with ${{A}^{\omega,{n}}}$, and ${(B^{m})}^{\omega}$ with ${{B}^{\omega,{m}}}$. If $A$ ($B$) is a field then $A^{n}$ ($B^{m}$) can be made into a field in the usual way.
Example 4.2.
Stream circuits (Rutten, 2008) are clocked, hardware-like finite structures, which are obtained by (finite) compositions of the (rational) stream transformers of the form ($r\in\mathbb{R}$)
@@
for multiplying a stream by a constant, adding two streams, copying, and delaying a stream. Such a stream circuit is visualized in Figure 1.
[…]
Figure 1. Finite Stream Circuit.
Example 4.3.
@@
for the output stream $y$ to obtain $y=(\nicefrac{{1}}{{1-X}})\cdot z$. Thus, the (rational) stream transformer ${\mathbb{R}}^{\omega}\to{\mathbb{R}}^{\omega}$ that is implemented by the feedback circuit in Figure 1 is given by $z\mapsto(\nicefrac{{1}}{{1-X}})\cdot z$. Since $\nicefrac{{1}}{{1-X}}=(1,1,1,\ldots)$, the output stream $y$ is, by stream multiplication (8), of the form $(\sum_{k=0}^{n}[X^{k}]z)_{n\in\mathbb{N}}$.
Composing Transformers.
@@
Example 4.4 (Functional Composition).
A formal power series may also be viewed as a stream transformer $S(X)$, where $X$ is now instantiated with an input stream, say, $T(X)$ by the ”composition” $S(T(X))$. Formally, the functional composition $T\,\circ\,S$ of two power series $S(X)\,\coloneqq\,\sum_{k\geq 0}a_{k}X^{k}$ and $T(X)\,\coloneqq\,\sum_{k\geq 1}b_{k}X^{k}$, that is, $b_{0}=0$, is a power series with coefficients
(14) $c_{k}$ $\,\coloneqq\,\sum_{n\geq 0}\sum_{j_{1}+\ldots+j_{n}=k}b_{n}a_{j_{1}}\ldots a_{j_{n}}\mbox{\@.}$
Inverses for the functional composition (14) of two streams exist under certain conditions, and may be used for reverse computation.
Definition 4.5.
@@
$S;T$ is sequential composition, $S\otimes T$ is parallel composition, $S\mathrm{\sqcup}T$ realizes angelical and $S\mathrm{\sqcap}T$ demonic nondeterminism, which is possibly unbounded.
These definitions work for arbitrary vector- and multivalued maps $S$, $T$ for which the composition $T(S(f))$ is well-defined (cmp. Example 4.4).
Refinement.
@@
(15) $S\preceq T$ $\,\coloneqq\,(\forall{f\in{A}^{\omega}})\,{S(f)\subseteq T(f)}\mbox{\@,}$